OpenStudy (anonymous):

Write an equation of the line with an X-Intercept of (5,0) and a Y-Intercept of (0,15)

OpenStudy (mrnood):

do you know the equation for a straight line y= mx + b ? (the value for m is the slope and the value for b is the y intercept)

OpenStudy (kkutie7):

y2-y1/x2-x1 is how you find the slope in the equation above m=slope. x is left alone. use this equation instead y=m(x-x1)+y1 goodluck =D